# Adversarial KA

### v1

@@ Abstract.
Regarding the representation theorem of Kolmogorov and Arnold (KA) as an algorithm for representing or <<expressing>> functions, we test its robustness by analyzing its ability to withstand adversarial attacks. We find KA to be robust to countable collections of continuous adversaries, but unearth a question about the equi-continuity of the outer functions that, so far, obstructs taking limits and defeating continuous groups of adversaries. This question on the regularity of the outer functions is relevant to the debate over the applicability of KA to the general theory of NNs.
@@ 1. Introduction
[…]
@@
Much debate ensued [6, 7] over the practical implications of KA for the central questions of trainability and generalizability. While the early consensus seemed to be that KA was not of practical value in the evolving technology, later authors [8, 9] have had a more optimistic view, focusing on the philosophy of the KA proof which hints at an optimal division of labor between the shallow and deep layers on a NN: data preparation / dynamics applied to the data. KA gives a <<universal>> presentation of the data on a single hidden layer and then builds a dynamic to converge to the final layer functions ($g$) needed to represent a desired multivariate ($f$).
@@
KA is essentially an algorithm for building a special type (shallow, one hidden layer; but highly non-linear) of NN to represent a given continuous function. In computer science, coding theory, and the theory of error correction, the ability of an algorithm to withstand adversarial noise serves as a fundamental measure of its robustness and power. This is typically the most demanding test to which algorithms are subjected. In this spirit, we investigate KA to see how much <<untapped>> expressivity still resides within the basic proof method. In our context, the adversary acts on the hidden layer and its effects must be countered by a new choice of $g$. Unlike error correcting codes, we presume knowledge of exactly how the adversary acted. This does not make the task of countering the adversary trivial, because we are not allowed to reach back and invert its effect on the hidden layer; we are only allowed to adjust the outer functions $g$.
@@
We note that in a different context the theory of NNs has already been much influenced by efforts to find and tame adversarial examples [10]. For example, NNs which easily distinguish real world examples of cats from dogs but can often be spoofed by carefully engineered pictures that look to humans exactly like <<cat>> or <<dog>> but that the NN will misclassify. Our adversary is somewhat different; it is not acting on inputs but rather corrupting the inner workings of the NN.
@@
Our conclusion is a bit surprising. We find in rather general circumstances that adversaries drawn from countable collections of homeomorphisms can be defeated, showing remarkable robustness. But we unearth a novel question of equi-continuity of our final functions $g$ (equi- in the adversary parameter), which we have been unable to answer. A positive answer means KA can defeat adversaries drawn from many continuous groups. So, KA’s exact level of adversarial robustness is framed as our final Questions 3.4, 3.5.
@@ 2. Background and statements
[…]
We consider only continuous functions with sup-norms, so we will omit it from the statements and definitions. Sometimes we will use the notation $C(X,Y)$ for such spaces of functions $X\to Y$. For a function $f\colon X\to Y\subset\mathbb{R}^{d}$ we denote by $f_{i}\colon X\to\mathbb{R}$ the composition of $f$ and projection on the $i$th coordinate.
The functions $I\to I^{2n+1}$ we will call tuples, since they may be considered as the tuples of $12n+$ 1 functions. For such functions we consider the maximum norm.
@@
In other words, Theorem 2.1 states that we may overcome an adversary acting as in Figure 1. Note that we cannot overcome an adversary acting as in Figure 2, since such general homeomorphism could undo all the preparations of the first, <<blocky>> layer map and render the composition $\mathbb{R}^{n}\to\mathbb{R}^{2n+1}\to\mathbb{R}^{2n+1}$ merely an inclusion into the first $n$ coordinates of $\mathbb{R}^{2n+1}$. For a specific example, one may take $n=2$ and $f(x,y)=xy$. It is easy to check that $f$ cannot be reconstructed at the four points (±$(\pm 1,\pm 1)$ even as $f(x,y)=g_{1}(x)+g_{2}(y)$. The difference between cases is that in the first case the hidden neurons receive information, it is <<scrambled>>, and then added, while in the second case the information is added and then scrambled.
Theorem 2.1 covers several interesting countable groups of adversaries. For example, $\mathcal{H}$ may consist of affine transformations defined by matrices and vectors with rational coefficients, or the affine group over number field, $AGL_{n}(\mathbb{K})$, or the group generated by homeomorphisms whose coordinate functions are polynomials over some number field.
[…]
Figure 3. Commuting of the <<extension>> from one to two variables and the affine transform $A$
@@
i.e. affine adversaries commute (up to a translation) with <<extension>> from one to more ($n$) variables (see Figure 3 where it is shown explicitly for $n=2$). Theorem 2.1 implies that we may overcome an affine adversary acting on all input neurons at once, as in Figure 2. The latter does not hold for arbitrary adversaries, as we explained right after Theorem 2.1.
Note that there is one more natural way to place an adversary: it may act before the summation, but after the multiplication, i.e. the approximation would look like
@@
Obviously, for an affine adversary this is just the same function (up to a translation). For a general case, however, we do not know how to overcome such adversaries.
If there are several adversaries, then the approximation is still possible in spite of the following result.
@@
Since $g^{f,h}$ are uniformly continuous and (1) is an open condition, it is sufficient to prove the denseness. That is, for any $\phi^{0}\colon I\to I^{5}$ we must construct $\phi\colon I\to I^{5}$ such that $\left\lVert\phi-\phi^{0}\right\rVert<\varepsilon$ and (1) holds for some appropriate $g^{f,h}$.
@@
It remains to construct $g^{f,h}$. A red rectangle of rank $i$, $i=1,\ldots,5$, in $I^{2}$, is defined to be the Cartesian product of any two red intervals of rank $i$. Denote these rectangles by $R_{i,1},\ldots,R_{i,r}$ for some $r$ (probably dependent of $i$) in any order [^3] In the case of general $n\geqslant 2$ they would be red $n$-rectangular solids..
It can be easily shown that any $x\in I$ lies in red intervals of at least four different ranks. So, any $(x,y)\in I^{2}$ lies in red rectangles of at least three different ranks [^4] In the case $n\geqslant 2$ any point from $I^{n}$ lies in red $n$-rectangular solids of at least $n+1$ different ranks.. On such rectangles $R_{j,r}$
@@
Finally, define a piecewise linear $g^{f,h}\colon\mathbb{R}\to\mathbb{R}$ by the <<forced>> values $g^{f,h}(\Phi^{h}_{j,r}):=\frac{1}{3}f(\ell_{j,r})$ for each red rectangle $R_{j,r}$. This definition makes sense since all $\Phi_{j,r}^{h}$ are distinct, because $\sqrt{2}$ is irrational.
Clearly it is possible to extend $g^{f,h}$ from its <<forced>> values to ensure that $\left\lVert g^{f,h}\right\rVert\leqslant\frac{1}{3}$.
@@
$\left\lvert f(x,y)-\sum_{j=1}^{5}g\bigl{(}\Phi_{j}^{h}(x,y)\bigr{)}\right\rvert<\alpha+\frac{2}{3}\quad\text{for any}\quad(x,y)\in I^{2}.$
Take any $x,y\in I$. As we already know, there are at least three red rectangles $R_{j_{s},r_{s}}$, $s\in\{1,2,3\}$, containing the point $(x,y)$. For them we have $g^{f,A}(\Phi_{j_{s},r_{s}}^{h})=\frac{1}{3}f(\ell_{j_{s},r_{s}})$. These values differ from $\frac{1}{3}f(x,y)$ by less than $\frac{\alpha}{3}$. Hence
@@
Let $n>1$ be an integer and $\lambda\in(\frac{2n+1}{2n+2},1)$. Let =$\mathcal{H}=\{h_{q}\}$ be a countable set of homeomorphisms $\mathbb{R}^{2n+1}\to\mathbb{R}^{2n+1}$. There exists a fixed $\phi\colon I\to I^{2n+1}$ so that given any $f\colon I^{n}\to\mathbb{R}$ and any $h\in\mathcal{H}$, there is uniformly continuous $g\colon\mathbb{R}\to\mathbb{R}$ with $\left\lVert g\right\rVert\leqslant\frac{1}{n+1}\left\lVert f\right\rVert$ satisfying
@@
(In fact, this intersection is a dense set of <<2nd category>>.) Choose $\phi$ from this intersection. For this $\phi$ choose maps $g^{f_{p},h_{q}}$ from the definition of $U_{f_{p},h_{q}}$.
@@
Since $\left\lVert g_{m}\right\rVert\leqslant\frac{1}{n+1}\left\lVert f-\sum\limits_{k=0}^{m-1}\operatorname{Approx}(g_{k},h)\right\rVert\leqslant\frac{1}{n+1}\lambda^{m-1}\left\lVert f\right\rVert$, the functional series $\sum\limits_{m=0}^{\infty}g_{m}$ converges uniformly on $\mathbb{R}$ to a function $g\colon\mathbb{R}\to\mathbb{R}$. Then
@@
Now we are going to give the proof of Theorem 2.2. The proof is analogous to the proof of Theorem 2.1, so we give only the sketch.
@@
(2) $\left\lvert f(x_{1},\ldots,x_{n})-\sum_{j=1}^{2n+1}g^{f,\{h^{i}\}}\Bigl{(}h^{1}_{j}\bigl{(}\phi^{1}(x_{1})\bigr{)}+\ldots+h^{n}_{j}\bigl{(}\phi^{n}(x_{n})\bigr{)}\Bigr{)}\right\rvert<\frac{n}{n+1}+\alpha\qquad\text{for any}\quad x_{1},\ldots,x_{n}\in I.$
@@
Since $g^{f,\{h^{i}\}}$ are uniformly continuous and (2) is an open condition, it is sufficient to prove the denseness. That is, for any $\phi^{0,1},\ldots,\phi^{0,n}\colon I\to I^{5}$ we must construct $\phi^{1},\ldots,\phi^{n}\colon I\to I^{5}$ such that $\left\lVert\phi^{i}-\phi^{0,i}\right\rVert<\varepsilon$ for each $i$ and (2) holds for some appropriate $g^{f,\{h^{i}\}}$.
The proof differs from the proof of Lemma 3.1 only in the construction of functions $\phi^{i}$. Take red intervals from the proof of Lemma 3.1. Take $\gamma_{i}$ to be square roots of distinct prime numbers. Now we specify $\phi^{i}$ somewhat indirectly by giving the following conditions on $\phi^{h^{i}}:=h^{i}\phi^{i}$. Since $h^{i}$ are invertible this suffices.
@@
    $\phi^{h^{i}}$ has range in $h^{i}I^{5}$ for each $i=1,\ldots,n$;
@@
Then denote

### v3

# Kolmogorov–Arnold stability

@@ Abstract.
Regarding the representation theorem of Kolmogorov and Arnold (KA) as an algorithm for representing or <<expressing>> functions, we test its robustness by analyzing its stability to withstand re-parameterizations of the hidden space. One may think of such re-parameterizations as the work of an adversary attempting to foil the construction of the KA outer function. We find KA to be stable under countable collections of continuous re-parameterizations, but unearth a question about the equi-continuity of the outer functions that, so far, obstructs taking limits and defeating continuous groups of re-parameterizations. This question on the regularity of the outer functions is relevant to the debate over the applicability of KA to the general theory of NNs.
S. V. Dzhenzher: Moscow Institute of Physics and Technology, Dolgoprudny 141701, Russia. sdjenjer@yandex.ru
M. H. Freedman: (Senior Research Scientist) CMSA-Harvard, and Google Research. mikehartleyfreedman@outlook.com
@@ 1. Introduction
Hilbert’s 13th problem [21] concerns the structure of expressions for the local motion of a root as the coefficients of a polynomial vary. His starting point was the quadratic formula known to all school children: $x=\frac{-b\pm\sqrt{b^{2}-4ac}}{2a}$ and its generalization to equations of degree $3$ and $4$. In these formulas $x$, a function of several variables, is built as a composition of functions in which the only functions with multiple inputs are ++ and $\cdot$. In fact, exploiting the single variable functions $\log$ and $\exp$, $\cdot$ becomes redundant. So Hilbert asked, in modern language, about the expressivity of what would later be called a Neural Network (NN). He proposed that mathematicians devise vast generalizations of Abel’s theorem on quintics, and demonstrate the limitations of compositions where the only multivariate function is ++. Perhaps inadvisably [^1] If we may be so bold as to critique Hilbert, his ‘error’ was underestimating the ability of continuous functions to confound dimensional arguments, a prime example being the space filling map $h\colon[\,0,1\,]\to[\,0,1\,]^{2}$ defined by $h(0.x_{1}x_{2}\ldots):=(0.x_{1}x_{3}\ldots,0.x_{2}x_{4}\ldots)$, i.e. the coordinate splitting map., Hilbert permitted merely continuous functions within the allowed expressions and thus opened the door for Kolmogorov and Arnold (KA) to refute (a part of) Hilbert’s intuition with their famous twin papers [1, 11]. It took 30 years [8, 19, 15] for the computer science community to realize KA had already addressed their most basic question: expressivity.
@@
Much debate ensued [6, 12] over the practical implications of KA for the central questions of trainability and generalizability. While the early consensus seemed to be that KA was not of practical value in the evolving technology, later authors [13, 4] have had a more optimistic view, focusing on the philosophy of the KA proof which hints at an optimal division of labor between the shallow and deep layers on a NN: data preparation / dynamics applied to the data. KA gives a ‘universal’ presentation of the data on a single hidden layer and then builds a dynamic to converge to the final layer functions ($g$) needed to represent a desired multivariate ($f$).
In this paper, we investigate the sensitivity and stability of KA NNs under continuous re-parameterizations of the hidden layer. More precisely, we are allowed to adjust the outer function $g$, but we are not allowed to reach back and invert the re-parameterization effects on the hidden layer.
@@
A metaphor for stability that we found useful, was to think of the intervening homeomorphism as the action of an adversary which can be overcome by a modification of the outer map(s). KA is essentially an algorithm for building a special type (shallow, one hidden layer; but highly non-linear) of NN to represent a given continuous function. In computer science, coding theory, and the theory of error correction, the ability of an algorithm to withstand adversarial noise serves as a fundamental measure of its robustness and power. This is typically the most demanding test to which algorithms are subjected. In this spirit, we investigate KA to see how much ‘untapped’ expressivity still resides within the basic proof method. In our context, the adversary acts on the hidden layer and its effects must be countered by a new choice of $g$. Unlike error correcting codes, we presume knowledge of exactly how the adversary acted.
@@
We note that in a different context the theory of NNs has already been much influenced by efforts to find and tame adversarial examples [7]. For example, NNs which easily distinguish real world examples of cats from dogs but can often be spoofed by carefully engineered pictures that look to humans exactly like ‘cat’ or ‘dog’ but that the NN will misclassify. Our adversary is somewhat different; it is not acting on inputs but rather corrupting the inner workings of the NN.
@@
Our conclusion is a bit surprising. We find in rather general circumstances that KA persists under countable collections of re-parameterization homeomorphisms, but becomes obstructed for continuous collections. The main obstruction is the failure of equi-continuity of our final functions $g$ over the collection of re-parameterizations. On this obstruction we are unable to answer. A positive answer means KA can defeat re-parameterizations drawn from many continuous groups. So, KA’s exact level of stability is framed as our final Questions 5.1, 5.2.
@@ 2. Background and statements
Denote I:=[ 0,1]I:=[\,0,1\,].
We consider only continuous functions with sup-norms, so we will omit it from the statements and definitions. Sometimes we will use the notation $C(X,Y)$ for such spaces of functions $X\to Y$. For a function $f\colon X\to Y\subset\mathbb{R}^{d}$ we denote by $f_{i}\colon X\to\mathbb{R}$ the composition of $f$ and the projection on the $i$-th coordinate.
The functions $I\to I^{2n+1}$ we will call tuples, since they may be considered as the tuples of $2n+1$ functions. For such functions we consider the maximum norm.
@@
In other words, Theorem 2.1 states that we may overcome a re-parameterization (or an adversary) acting as in Figure 1. Note that we cannot overcome a re-parameterization acting as in Figure 2, since such general homeomorphism could undo all the preparations of the first, ‘blocky’ layer map and render the composition $\mathbb{R}^{n}\to\mathbb{R}^{2n+1}\to\mathbb{R}^{2n+1}$ merely an inclusion into the first $n$ coordinates of $\mathbb{R}^{2n+1}$. For a specific example, one may take $n=2$ and $f(x,y)=xy$. It is easy to check that $f$ cannot be reconstructed at the four points $(\pm 1,\pm 1)$ even as $f(x,y)=g_{1}(x)+g_{2}(y)$. The difference between cases is that in the first case the hidden neurons receive information, it is ‘scrambled’, and then added, while in the second case the information is added and then scrambled.
Theorem 2.1 covers several interesting countable groups of re-parameterizations. For example, $\mathcal{H}$ may consist of affine transformations defined by matrices and vectors with rational coefficients, or the affine group over number field, $AGL_{n}(\mathbb{K})$, or the group generated by homeomorphisms whose coordinate functions are polynomials over some number field.
$\phi_{i}(x)$$\phi(x)+\sqrt{2}\phi(y)$$A^{i}_{j}(\phi_{i}(x)+a_{i})$$A^{i}_{j}(\phi_{i}(x)+a_{i})+\sqrt{2}A^{i}_{j}(\phi_{i}(y)+a_{i})$$A^{i}_{j}(\phi_{i}(x)+\sqrt{2}\phi_{i}(y)+a_{i})$extensionAAextensionAAdiffer by $A^{i}_{j}(\sqrt{2}a_{i})$
Figure 3. Commuting of the ‘extension’ from one to two variables and the affine transform AA
@@
i.e. affine re-parameterizations commute (up to a translation) with ‘extension’ from one to more ($n$) variables (see Figure 3 where it is shown explicitly for $n=2$). Theorem 2.1 implies that we may overcome an affine re-parameterization acting on all input neurons at once, as in Figure 2. The latter does not hold for arbitrary re-parameterizations, as we explained right after Theorem 2.1.
Note that there is one more natural way to place a re-parameterization: it may act before the summation, but after the multiplication, i.e. the approximation would look like
@@
Obviously, for an affine re-parameterization this is just the same function (up to a translation). For a general case, however, we do not know how to overcome such re-parameterizations.
If there are several re-parameterizations, each acting on a different inner function, then the approximation is still possible in spite of the following result.
@@
First, let us demonstrate the openness of $U_{f,h}$. Suppose that $\phi\in U_{f,h}$. Take $g=g^{f,h}$ from the definition of $U_{f,h}$. Set
$\varepsilon:=\frac{2}{3}+\alpha-\left\lVert f-\operatorname{Approx}(g,h)\right\rVert>0.$
By the uniform continuity of $g$, there is $\delta>0$ such that $\left\lvert g(x)-g(y)\right\rvert<\frac{\varepsilon}{5}$ as $\left\lvert x-y\right\rvert<\delta$. Since on the cube $I^{5}$ the homeomorphism $h$ is uniformly continuous, there is $\omega>0$ such that $\left\lVert h(x)-h(y)\right\rVert<\frac{\delta}{1+\sqrt{2}}$ as $\left\lVert x-y\right\rVert<\omega$. Now we show that any $\psi$ which is $\omega$-close to $\phi$ also lies in $U_{f,h}$. Fix such $\psi$. Then for any $j=1,\ldots,5$ and $x,y\in I$ we have
$\left\lvert h_{j}\bigl(\phi(x)\bigr)+\sqrt{2}h_{j}\bigl(\phi(y)\bigr)-h_{j}\bigl(\psi(x)\bigr)-\sqrt{2}h_{j}\bigl(\psi(y)\bigr)\right\rvert<\left(1+\sqrt{2}\right)\frac{\delta}{1+\sqrt{2}}=\delta.$
For the convenience, till the end of this paragraph let us denote by $\operatorname{Approx}_{\phi}$ and $\operatorname{Approx}_{\psi}$ the approximator constructed by $\phi$ and $\psi$, respectively. Hence
$\left\lvert\operatorname{Approx}_{\phi}(g,h)-\operatorname{Approx}_{\psi}(g,h)\right\rvert<5\cdot\frac{\varepsilon}{5}=\varepsilon.$
Finally,
$\left\lvert f-\operatorname{Approx}_{\psi}(g,h)\right\rvert<\left\lvert f-\operatorname{Approx}_{\phi}(g,h)\right\rvert+\varepsilon=\frac{2}{3}+\alpha,$
which proves the openness.
In the rest part of the proof we prove the denseness of $U_{f,h}$. That is, for any $\phi^{0}\colon I\to I^{5}$ we must construct $\phi\colon I\to I^{5}$ such that $\left\lVert\phi-\phi^{0}\right\rVert<\varepsilon$ and (1) holds for some appropriate $g^{f,h}$. The idea of the proof is to change the coordinate system on the one twisted by the re-parametrization $h$, and then to repeat the classical proof.
@@
It remains to construct $g^{f,h}$. A red rectangle of rank $i$, $i=1,\ldots,5$, in $I^{2}$, is defined to be the Cartesian product of any two red intervals of rank $i$. Denote these rectangles by $R_{i,1},\ldots,R_{i,r}$ for some $r$ (probably dependent on $i$) in any order [^3] In the case of general $n\geqslant 2$ they would be red $n$-rectangular solids..
In this paragraph we give the well-known argument in order to show that any $x\in I$ lies in red intervals of at least four different ranks; see for example [2, Fundamental Lemma 1]. The red ractangles of rank $i$ are of the form
$I\cap\left[\,\frac{5r+i-4}{N},\frac{5r+i}{N}\,\right].$
Then
$x\in I\cap\left[\,\frac{5r+i-4}{N},\frac{5r+i}{N}\,\right]\Longleftrightarrow Nx\in\left[\,5r+i-4,5r+i\,\right].$
Taking $r$ such that $Nx\in[\,5r-4,5r+1)$, we obtain that at least for four different $i$ the inclusion $Nx\in\left[\,5r+i-4,5r+i\,\right]$ holds.
Hence, any $(x,y)\in I^{2}$ lies in red rectangles of at least three different ranks [^4] In the case $n\geqslant 2$ any point from $I^{n}$ lies in red $n$-rectangular solids of at least $n+1$ different ranks.. On such red rectangles $R_{j,r}$
@@
Finally, define a piecewise linear $g^{f,h}\colon\mathbb{R}\to\mathbb{R}$ by the ‘forced’ values $g^{f,h}(\Phi^{h}_{j,r}):=\frac{1}{3}f(\ell_{j,r})$ for each red rectangle $R_{j,r}$. This definition makes sense since all $\Phi_{j,r}^{h}$ are distinct. Indeed, on two different rectangles $R_{j,r}$ and $R_{j^{\prime},r^{\prime}}$ we have
$\Phi_{j,r}^{h}=p+\sqrt{2}q\quad\text{and}\quad\Phi_{j^{\prime},r^{\prime}}^{h}=p^{\prime}+\sqrt{2}q^{\prime},$
where $p,p^{\prime},q,q^{\prime}\in\mathbb{Q}$ and at least one of the inequalities $p\neq p^{\prime}$, $q\neq q^{\prime}$ holds, and that is why $\Phi_{j,r}^{h}\neq\Phi_{j^{\prime},r^{\prime}}^{h}$.
Clearly, it is possible to extend $g^{f,h}$ from its ‘forced’ values to ensure that $\left\lVert g^{f,h}\right\rVert\leqslant\frac{1}{3}$. For that, one just needs to extend $g^{f,h}$ piecewise linearly between the leftmost and the rightmost points $\Phi^{h}_{j,r}$, and, on the two remaining rays, make it remain constant.
@@
$\left\lvert f(x,y)-\sum_{j=1}^{5}g^{f,h}\bigl(\Phi_{j}^{h}(x,y)\bigr)\right\rvert<\alpha+\frac{2}{3}\quad\text{for any}\quad(x,y)\in I^{2}.$
Take any $x,y\in I$. As we already know, there are at least three red rectangles $R_{j_{s},r_{s}}$, $s\in\{1,2,3\}$, containing the point $(x,y)$. For them we have $g^{f,h}(\Phi_{j_{s},r_{s}}^{h})=\frac{1}{3}f(\ell_{j_{s},r_{s}})$. These values differ from $\frac{1}{3}f(x,y)$ by less than $\frac{\alpha}{3}$. Hence
@@
Let $n>1$ be an integer and $\lambda\in(\frac{2n+1}{2n+2},1)$. Let $\mathcal{H}=\{h_{q}\}$ be a countable set of homeomorphisms $\mathbb{R}^{2n+1}\to\mathbb{R}^{2n+1}$. There exists a fixed $\phi\colon I\to I^{2n+1}$ such that given any $f\colon I^{n}\to\mathbb{R}$ and any $h\in\mathcal{H}$, there is a uniformly continuous $g\colon\mathbb{R}\to\mathbb{R}$ with $\left\lVert g\right\rVert\leqslant\frac{1}{n+1}\left\lVert f\right\rVert$ satisfying
@@
(In fact, this intersection is a dense set of ‘2nd category’.) Choose $\phi$ from this intersection. For this $\phi$ choose maps $g^{f_{p},h_{q}}$ from the definition of $U_{f_{p},h_{q}}$.
@@
Since $\left\lVert g_{m}\right\rVert\leqslant\frac{1}{n+1}\left\lVert f-\sum\limits_{k=0}^{m-1}\operatorname{Approx}(g_{k},h)\right\rVert\leqslant\frac{1}{n+1}\lambda^{m-1}\left\lVert f\right\rVert$, the functional series $\sum\limits_{m=0}^{\infty}g_{m}$ converges uniformly on $\mathbb{R}$ to a uniformly continuous function $g\colon\mathbb{R}\to\mathbb{R}$. Then for any $j=1,\ldots,2n+1$
$\sum_{m=0}^{\infty}g_{m}\left(\gamma_{1}t_{j,1}+\ldots+\gamma_{n}t_{j,n}\right)=g\left(\gamma_{1}t_{j,1}+\ldots+\gamma_{n}t_{j,n}\right),$
where the convergence is uniform for any $t_{j,i}\in\mathbb{R}$. Finally,
@@
4. Proof of Theorem 2.2
In this section we give the proof of Theorem 2.2. Since the proof is analogous to the proof of Theorem 2.1, the reader may skip it and go right to §5.
For fixed tuples $\phi^{1},\ldots,\phi^{n}\colon I\to I^{2n+1}$, a function $g\colon\mathbb{R}\to\mathbb{R}$, and homeomorphisms $h^{1},\ldots,h^{n}\colon\mathbb{R}^{2n+1}\to\mathbb{R}^{2n+1}$ define the approximator $\operatorname{Approx}(g,\{h^{i}\})\colon I^{n}\to\mathbb{R}$ by
$\operatorname{Approx}(g,\{h^{i}\})(x_{1},\ldots,x_{n}):=\sum_{j=1}^{2n+1}g\Bigl(h^{1}_{j}\bigl(\phi^{1}(x_{1})\bigr)+\ldots+h^{n}_{j}\bigl(\phi^{n}(x_{n})\bigr)\Bigr).$
@@
(2) $\left\lvert f(x_{1},\ldots,x_{n})-\operatorname{Approx}(g^{f,\{h^{i}\}},\{h^{i}\})\right\rvert<\frac{n}{n+1}+\alpha$
for any $x_{1},\ldots,x_{n}\in I$.
@@
First, let us demonstrate the openness of $U_{f,h}$. Suppose that $\phi\in U_{f,\{h^{i}\}}$. Take $g=g^{f,\{h^{i}\}}$ from the definition of $U_{f,\{h^{i}\}}$. Set
$\varepsilon:=\frac{n}{n+1}+\alpha-\left\lVert f-\operatorname{Approx}(g,\{h^{i}\})\right\rVert>0.$
By the uniform continuity of $g$, there exists $\delta>0$ such that $\left\lvert g(x)-g(y)\right\rvert<\frac{\varepsilon}{2n+1}$ as $\left\lvert x-y\right\rvert<\delta$. Since on the cube $I^{2n+1}$ the homeomorphisms $h^{i}$ are uniformly continuous, there is common $\omega>0$ such that $\left\lVert h^{i}(x)-h^{i}(y)\right\rVert<\frac{\delta}{n}$ as $\left\lVert x-y\right\rVert<\omega$. Now we show that any $\psi$ which is $\omega$-close to $\phi$ also lies in $U_{f,\{h^{i}\}}$. Fix such $\psi$. Then for any $j=1,\ldots,2n+1$ and $x_{1},\ldots,x_{n}\in I$ we have
$\left\lvert h^{1}_{j}\bigl(\phi^{1}(x_{1})\bigr)+\ldots+h^{n}_{j}\bigl(\phi^{n}(x_{n})\bigr)-h^{1}_{j}\bigl(\psi^{1}(x_{1})\bigr)-\ldots-h^{n}_{j}\bigl(\psi^{n}(x_{n})\bigr)\right\rvert<n\frac{\delta}{n}=\delta.$
For the convenience, till the end of this paragraph let us denote by $\operatorname{Approx}_{\phi}$ and $\operatorname{Approx}_{\psi}$ the approximator constructed by $\phi$ and $\psi$, respectively. Hence
$\left\lvert\operatorname{Approx}_{\phi}(g,h)-\operatorname{Approx}_{\psi}(g,h)\right\rvert<(2n+1)\frac{\varepsilon}{2n+1}=\varepsilon.$
Finally,
$\left\lvert f-\operatorname{Approx}_{\psi}(g,h)\right\rvert<\left\lvert f-\operatorname{Approx}_{\phi}(g,h)\right\rvert+\varepsilon=\frac{n}{n+1}+\alpha,$
which proves the openness.
In the rest part of the proof we prove the denseness. That is, for any $\phi^{0,1},\ldots,\phi^{0,n}\colon I\to I^{2n+1}$ we must construct $\phi^{1},\ldots,\phi^{n}\colon I\to I^{2n+1}$ such that $\left\lVert\phi^{i}-\phi^{0,i}\right\rVert<\varepsilon$ for each $i$ and (2) holds for some appropriate $g^{f,\{h^{i}\}}$.
The proof differs from the proof of Lemma 3.1 only in the construction of functions $\phi^{i}$. Take red intervals of rank $i\equiv 0,\ldots,2n\pmod{2n+1}$ from the proof of Lemma 3.1. Take $\gamma_{i}$ to be square roots of pairwise distinct prime numbers. Now we specify $\phi^{i}$ somewhat indirectly by giving the following conditions on $\phi^{h^{i}}:=h^{i}\phi^{i}$. Since $h^{i}$ are invertible this suffices.
@@
    $\phi^{h^{i}}$ has range in $h^{i}I^{2n+1}$ for each $i=1,\ldots,n$;
@@
Now taking red $n$-rectangular solids of rank $i$ as the Cartesian products of red intervals of rank $i$, we observe that any point $(x_{1},\ldots,x_{n})\in I^{n}$ lies in red $n$-rectangular solids of at least $n+1$ different ranks. On such $n$-rectangular solids $R_{j,r}$